The asolute abundance ratio of N14/N15 in atmospheric nitrogen found to be 272.  Calculate atomic weight of atomespheric Nitrogen?

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi kiran,

N14/N15=272/1

let a sample contains 1 atom of N15 then it contains 272 N14 atoms...

Atomic weight= (272*14 + 1* 15)/273

At. Wt.=14.0037